1)   Traffic lights with a turning arrow are red 60% of the time. If you approach 5 lights in a row, find the probability of stopping:

a.     Exactly three times.

 

 

 

 

b.    At least two times.

 


1
Expert's answer
2021-05-17T00:55:02-0400

Solution:

Given, "p=60\%=0.6, q=0.4,=5"

"X\sim Bin(n,p)"

(a) "P(X=3)=^5C_3 (0.6)^3(0.4)^2=0.3456"

(b) "P(X\ge2)=1-P(X<2)=1-[P(X=0+P(X=1)]"

"=1-[^5C_0 (0.6)^0(0.4)^5+^5C_1 (0.6)^1(0.4)^4]
\\=0.91296"